November 19, 2014 Calibration
GlueX Calibration Meeting
Wednesday, October 22, 2014
11:00 am, EDT
JLab: CEBAF Center, F326
Connection Using Bluejeans
- To join via Polycom room system go to the IP Address: 18.104.22.168 (bjn.vc) and enter the meeting ID: 630804895.
- To join via a Web Browser, go to the page  https://bluejeans.com/630804895.
- To join via phone, use one of the following numbers and the Conference ID: 630804895
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
- More information on connecting to bluejeans is available.
- Status Update
- Calibration status/updates
- Fall 2014 Commissioning Simulations updates
- Data Monitoring
Attending: Sean (NU); Kei, Simon, Lubomir, Mark I., Justin, Nathan S., Will McG., Naomi, Mike S. (JLab); Paul (CMU); Eric (FIU)
The target conditions are being tracked as a function of run, and in the process a problem where the run number wasn't always being saved in the EVIO files was found and fixed.
We then reviewed the status of various subdetectors.
We reviewed an email from Zisis. Mark D. extracted the attenuation lengths. Will M. is working on pi0 reconstruction. Tegan is looking at the thresholds in various runs. Andrei is extracting the fADC to physical unit calibration using cosmic data.
No pi0 peaks are found yet. The beam data studies are statistically limited.
Lubomir reported that last night they took several hours of data with the solenoid turned off. The data looked to be of good quality, with many reconstructed tracks. This data is allowing him to take a preliminary look at the alignment, e.g. the rotations of the different packages. Vlad has been looking at the cathode strip calibration using cosmic data, and he is starting to use the same technique with the new beam data. Preliminary results are encouraging.
The TDC readout for the wires is still a problem, which is limiting the obtainable time resultion, and hindering the calibration of the time-to-distance functions.
Simon has been analyzing the data and has posted several plots to the logbook. He is looking at matching tracks to other detectors. Matching to the start counter and FCAL is working well, and he has been able to image the start counter. For the FCAL, right now he can only match to hits not clusters. He is not able to match to the TOF hit positions.
Having a BCAL trigger would be useful for getting tracks that illuminate more of the FDC. Having tracks from the active target is useful as well. Simon's made several plots of track vertex location, and sees what looks like some tracks coming from active target.
10-20% of the tracks look like they come from the beam halo or interactions in the air.
Naomi and Mike said that they have taken data at three high voltage settings as part of a scan to find the best high voltage set point.
Simon has found that the quality of the reconstructed tracks with the field off is problematic and this is leading to problems matching to BCAL hits. The matching with the field on is fine. He is looking into this.
Sasha O. has been working on the TOF calibration primarily with cosmic data, and has been starting to analyze photon beam data. He's looked at run 995, but this did not have TDC information. Hopefully the data taken over the past night will be more useful. The fADC pulse integrals are also not understood.
Eric gave a comprehensive look at the start counter. The ADC and TDC readouts are now working well. There is progress on understanding the tilt of the SC, and it is expected to be fixed when mounted onto the hydrogen target. The relative timing of the SC TDCs with respect to BCAL seems to be okay, but the relationship to the FDC or TOF needs to be understood. This might be fixed with updated trigger firmware. See Mark Ito's talk in the online meeting this afternoon for more details. Work will soon shift to determining calibration constants with actual data.
Lubomir asked if Eric was calculating the TDC times in reference to the trigger signal. Eric said he was not since he didn't know how to get at this information. Lubomir said that he would email Eric about how to get that information.
New simulations will be generated corresponding to the current no B-field runs. The main change is that the beam profiler in the collimator cave was moved to in front of the photon beam dump.
For the detector rate analysis, one important missing components is understanding what the detector thresholds are in terms of physical units. Right now only the calorimeters know this. Mark D. reported that the BCAL threshold is around 10 MeV. Sean will follow up with Matt S. and the FCAL group to find their threshold.
Paul reported that the plan is to rerun the offline monitoring each Friday with the intent that the results should be on the website by the following Monday. The latest results are currently available on the web.
Everyone should look at these plots and give feedback for what would be useful for their detectors. During some later discussion, Mark I. suggested that this effort may need more advertisement. After some discussion, Kei agreed to give a brief, few-minute overview at the run coordination meeting tomorrow. Paul agreed to put some links to these pages on the front of the GlueX wiki.
The online monitoring output still has issues. Naomi and Mike S. reported problems using the RootSpy GUI on the online machines.
Kei gave a more detailed overview of the current state of this effort. This included an overview of the current runs, current issues with the data, the various webpages that are available, and an example comparison of different runs.
Sean reported that progress is being made on storing meta information in the database. The current plan is to store the official list of run conditions in the monitoring DB. Currently he is compiling data from several locations, and this list of runs is being used for the website display. There is ongoing work between him and David L. to store selected EPICS information in the data stream, to track this information in the offline framework, and to fill the run data table from the counting room. There is progress being made on this front, but it is being delayed by the DAQ issues over the past week.